y = x + 4
3y - 5x= 6 Solve by substitution

Found 2 solutions by scott8148, checkley77:
Answer by scott8148(6628)

substituting ___ 3(x + 4) - 5x = 6

3x + 12 - 5x = 6 ___ -2x = -6 ___ x = 3

substituting ___ y = 3 + 4 = 7

Answer by checkley77(12844)

y=x+4
3y-5x=6
3(x+4)-5x=6
3x+12-5x=6
-2x=6-12
-2x=-6
x=-6/-2
x=3 ans.
y=3+4
y=7 ans.
Proof:
3*7-5*3=6
21-15=6
6=6